mathslover (mathslover):

What is the value of (g) at north pole and at equator .. Is their any way we can find their value ?

mathslover (mathslover):

@Callisto

mathslover (mathslover):

can any1 help me ?

OpenStudy (anonymous):

Because the earth is very slightly flattened at the poles, so the North and South Poles are a bit closer to the center of the Earth than land on the Equator. And the South Pole is in Antarctica, which is a very high plateau, so the gravity there would be a bit less than that at the North Pole, which is in the middle of a deep ocean. The South Pole is colder than the North precisely because of this high elevation. g(at north pole)=GM/R^2=9.86m/s^2 g(at equator)= GM/R^2= 9.80m/s^2

OpenStudy (anonymous):

at North pole value of R = 6,356,800m at equator value of R= 6,378,100m

OpenStudy (anonymous):

@annas your explanation is correct. Earth bulges slightly in the middle, hence the variation in g. but the values i got are 9.83 at poles and 9.78 at equator. the mean of which comes out to be 9.8 which we generally use.

OpenStudy (anonymous):

the values i get are round about which were mentioned in my book.

OpenStudy (anonymous):

and this is the reason we use 9.8m/s^2 generally

mathslover (mathslover):

Oh thanks a lot @annas and @Vaidehi09 annas the value of R you got : if i didn't have this than i would not have been able to calculate that right ?

mathslover (mathslover):

@vishweshshrimali5 i want you to also enter here

OpenStudy (anonymous):

yes brother maths those values of R are necessary if you want to find values at poles and equator :) your welcome. i remember this because i got a question in my exams when i was in 10th grade.

mathslover (mathslover):

k thanks !! \[\huge{g_p>g_{eq.}}\] p--> poles and eq-->equator

OpenStudy (vishweshshrimali5):

@annas correct explanation @mathslover Actually, this deviation of value of acceleration of gravity forces us to take an average value instead of the correct value. Regarding your last reply, This can be the answer of a question like Why the gold measured at the equator increases its weight when brought to the poles ?

mathslover (mathslover):

Oh k ... thanks a lot @vishweshshrimali5 i got it perfectly now